Professor I know its a silly problem but I'm actually needing your help. Thanks!

A woman has $1.70 in dimes and nickels. She has 5 more dimes thank nickels. How many nickels does she have?

Answer by stanbon(57214)

A woman has $1.70 in dimes and nickels. She has 5 more dimes thank nickels. How many nickels does she have?
-----
Value Equation: 10d + 5n = 170 cents
Quantity Eq:::: d = n + 5
-------------------------------
Substitute for "d" and solve for "n":
10(n+5) + 5n = 170
10n + 50 + 5n = 170
15n = 120
n = 8 (# of nickels)
----
Solve for "d":
d = n+5
d = 8+5
d = 13 (# of dimes)
=======================
Cheers,
Stan H.
==================